OpenStudy (waheguru):

is it true when subtracting numbers if a positive number is being subtracted by negative number the answer will be positive no matter what and if the first nuber is negative the answer will be negative no matter what

OpenStudy (anonymous):

here are some examples: -2-4=-6, -2+4=2, 2-4=-2

OpenStudy (angela210793):

Examples: 1-(-3)=4>0 tht's always true -1-(-3)=-1+3=2 not always true

OpenStudy (waheguru):

so is my statement true or false

OpenStudy (anonymous):

when subtracting the answer takes the sign of the highest term

OpenStudy (dumbcow):

waheguru, when i said difference is always positive, its positive because of absolute value \[difference = |x_{2} - x{1}|\]

OpenStudy (waheguru):

im still confused

OpenStudy (waheguru):

so if u subtract a positive number from a negative the answer has to be potitive correct?

OpenStudy (dumbcow):

ok then ignore everything i said sorry to cause confusion just subtract final value minus initial value

OpenStudy (anonymous):

ok 6-(-4)= basically 6+4

OpenStudy (waheguru):

plz use simple tearms i ve onlygoing 2 grade 9

OpenStudy (anonymous):

6-4=2 because higher number is the 6 and its positive

OpenStudy (waheguru):

I WANNA KNOW IF MY STAMEMENT IS RIGHT OR WRONG

OpenStudy (anonymous):

it depends on the numbers if it is true or not

OpenStudy (waheguru):

can u give any examples

OpenStudy (anonymous):

ok 6-2 is a positive minus a negative do you agree

OpenStudy (waheguru):

isnt it a positive minus a positive

OpenStudy (anonymous):

6-13 is also a positive minus a negative however the minus is larger so the answer will be negative

OpenStudy (anonymous):

watch out there r exceptions. for ex, if u have 6--2, its actually 6+2

OpenStudy (waheguru):

jenni 6-2 is a posotove minus a positive a positive minus a negative would be 6-(-2)

OpenStudy (anonymous):

in the case of 6-2 yes it is positive

OpenStudy (anonymous):

The first part of your statement is true ex: 2-(-4)=6, 8-(-20)=28 the second part is false ex: -2-(-6)=4, -10-(-6)=-4

OpenStudy (angela210793):

I said as Mah.

OpenStudy (anonymous):

and yes there are cases called double negatives like 6--2 which is 6+2

OpenStudy (waheguru):

thx

OpenStudy (anonymous):

any time